OpenStudy (anonymous):

4(20-4p)-.75(4-16p) Simplify

OpenStudy (anonymous):

77.-4 p

OpenStudy (merchandize):

80-16p-300+1200p=0 -16p+1200p=300-80 1184p=220 p=220/1180 p=0.186